OpenStudy (buckybarnes):

The midpoint of UV is (5,-11) the coordinates of one endpoint are U(3,5). Find the coordinates of endpoint V.

OpenStudy (buckybarnes):

@3mar

OpenStudy (buckybarnes):

Would it be (7,-27)

OpenStudy (wholock):

First off, I love your username, second, the midpoint formula is (x1 + x2)/2 and (y1 + y2)/2 Since you have one point, just work it backwards. x = (x1 + x2)/2 5 = (3 + x)/2 <-- multiply both sides by 2 10 = 3 + x <-- Subtract 3 from both sides x = 7 and now for y, 11 = (5 + y)/ 2 <--- repeat steps, multiply by 2 22 = 5 + y <-- Subtract 5 17 = 7 So the other endpoint is (7, 17)

OpenStudy (3mar):

Your answer, sister, is correct.

OpenStudy (3mar):

(7,-27)

OpenStudy (3mar):

and check this for being sure.

OpenStudy (buckybarnes):

It's a -11 not 11 for the y part of the midpoint

OpenStudy (buckybarnes):

I did this -11 = (5 + y)/ 2 <--- multiply by 2 -22 = 5 + y <-- Subtract 5 -27=y

OpenStudy (wholock):

Okay, forgive me for missing the negative. Yes, you are correct!

OpenStudy (3mar):

Nice point, you are up!

OpenStudy (wholock):

Thanks for catching me @3mar

OpenStudy (buckybarnes):

Thank you guys so much

OpenStudy (3mar):

We are in the same place, brother! Be always at the top!

OpenStudy (3mar):

and thanks for the medal!